OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 4

jimthompson5910 (jim_thompson5910):

what do you get when you distribute 4(x+1)

OpenStudy (anonymous):

Distribute 4x+4=4x+4 0=0

OpenStudy (anonymous):

@jim_thompson5910 the answer is no solution or?

jimthompson5910 (jim_thompson5910):

supraron you got 0 = 0 is that equation true?

OpenStudy (anonymous):

the equation is true since x will always equal x but i dont think it asked to plug in a value for x

OpenStudy (anonymous):

it should just be x=x

OpenStudy (anonymous):

@jim_thompson5910 well it is true but since there are no numbers plugged it it cannot be a check so my guess is that the answer is no solution but not sure

jimthompson5910 (jim_thompson5910):

0 = 0 is true regardless of x what x is so 4(x + 1) = 4x + 4 is true for all values of x